# Likes, Budgets, and Equilibria: Designing Contests for Socially Optimal Advertising

### v1

@@ Abstract
Firms (businesses, service providers, entertainment organizations, political parties, etc.) advertise on social networks to draw people’s attention and improve their awareness of the brands of the firms. In all such cases, the competitive nature of their engagements gives rise to a game where the firms need to decide how to distribute their budget over the agents on a network to maximize their brand’s awareness. The firms (players) therefore need to optimize how much budget they should put on the vertices of the network so that the spread improves via direct (via advertisements or free promotional offers) and indirect marketing (words-of-mouth). We propose a two-timescale model of decisions where the communication between the vertices happen in a faster timescale and the strategy update of the firms happen in a slower timescale. We show that under fairly standard conditions, the best response dynamics of the firms converge to a pure strategy Nash equilibrium. However, such equilibria can be away from a socially optimal one. We provide a characterization of the contest success functions and provide examples for the designers of such contests (e.g., regulators, social network providers, etc.) such that the Nash equilibrium becomes unique and social welfare maximizing. Our experiments show that for realistic scenarios, such contest success functions perform fairly well.
@@ 1 Introduction
In a competitive market setting, potential customers (vertex in the social graph) have varying level of information available on different firms. In this paper, we measure this information disparity a vertex has about firms through its awareness level. It captures how prominently a firm occupies an agent’s attention or belief relative to its competitors. The awareness of the customers therefore consists of two components: (a) a direct marketing impact coming from the firms through advertisements, promotional offers, etc., and (b) an indirect impact coming from recommendations from their neighbors in the social graph. The consolidated impact on the awareness therefore depends on the network centralities of the vertex w.r.t. its neighbors and from the competing budget allocated to every node by all the firms. The competing budget component is modeled through the widely used contest success function (CSF) (Skaperdas, 1996). Our analysis considers that the firms play best response to each other’s committed budgets to reach equilibrium, and analyzes the welfare at such equilibria.
@@ 1.1 Our Contributions
The contributions of this paper is summarized as follows.
  • We introduce a fast and slow timescale model of awareness spread in a social network where on the firms update their budget strategies on the slow timescale and the individuals update their awareness on the faster timescale (Section˜2.1).
@@
  • On the slower timescale, we show that the budget strategy of the firms converge to a Nash equilibrium (NE) (Theorem˜1). However, the social welfare at the Nash equilibrium may not be optimal (˜1).
@@
Due to space limitations, some details of the proofs and experiments have been deferred to the supplementary material.
@@ 1.2 Related Work
Modeling and structural guarantees of competitive influence. A significant line of research has focused on formalizing competitive diffusion processes and analyzing the properties of equilibria. Foundational work by Bharathi et al. (2007) study the diffusion of competing innovations in social networks, proposing a tractable game-theoretic model, deriving a (11/e)(1-1/e)-approximation algorithm for best responses, and bounding the price of competition. Early works (Alon et al., 2010; Goyal and Kearns, 2012) formalize the game theoretic competitive influence maximizing model and emphasize the role of topology in the existence of Nash equilibrium and influence amplification. Zuo et al. (2022) introduce the Online Competitive Influence Maximization (OCIM) problem, and design bandit-based algorithms with sublinear regret. Tzoumas et al. Tzoumas et al. (2012) models the diffusion process of CIM in a simultaneous two-player non-cooperative game as a linear threshold model. Fazeli et al. Fazeli and Jadbabaie (2012a, b); Fazeli et al. (2016) develop a series of game-theoretic models for competitive product diffusion in social networks, where firms simultaneously allocate fixed budgets between seeding strategies and product quality, aiming to maximize the lower bound of their product adoptions.
@@ 2 Preliminaries
We consider a market with $m$ competing firms denoted by $\mathcal{M}=\{1,2,\ldots,m\}$, each having an equal and fixed marketing budget that they strategically allocate across users in a social network. Firms compete to attract the attention of these users by distributing their budgets to maximize brand visibility and influence. We consider awareness as a cardinal metric of a user’s attention towards a particular firm, which lies within $[0,1]$. This marketing competition is modeled as a finite non-cooperative game on a social network $\mathscr{G}=(\mathscr{N},\mathscr{E})$, where $\mathscr{N}=\{1,2,\ldots,n\}$ denotes the set of users in the network and $\mathscr{E}=(e_{ij})_{i,j\in\mathscr{N}}$ represents their influence structure through an adjacency matrix, with $e_{ii}=0$ for all $i\in\mathscr{N}$. Each firm aims to maximize the overall awareness of its brand across the network, subject to its budget constraints.
@@ 2.1 A Two-Timescale Communication Model
Figure 1 below illustrates the concept of communication flow on two-time scales: a faster one and a slower one.
[…]
Figure 1: The two-timescale concept
The slower timescale, $k={0,1,2,\ldots,\infty}$, governs firms’ budget allocations across each time period $k$ of length $T$, while the faster timescale, $t={0,T,2T,\ldots,\infty}$, captures how information spreads and stabilizes between users within each period. On the onset of the slower timescale, firms offer varying budget shares to different users (direct communication) with probability $(1-\alpha)$, who then disseminate the information received to their neighbors. This information propagation or awareness flow among the nodes (peer-to-peer communication) happens in the faster timescale with probability $\alpha$ and converges to its steady state value before the beginning of the next period of the slower timescale. We assume that $T$ is large such that the fast timescale process converges between two periods. This is often observed in practice where the advertisement budgets (and strategies) are updated by the firms once in every month or even more infrequently, while the viewers of the advertisements update their actions very frequently (perhaps every hour). In this paper, we will overload $t$ to be the running variable within two slower timescale periods $k$ and $k+1$, where $t$ runs from 0 to $T$.
@@ 2.1.1 Awareness levels of each firm
Each firm $s\in M$ chooses a strategy vector $\mathbf{b}_{s}=(b_{s1},\dots,b_{sn})$ representing the allocation of its marketing budget across the $n$ nodes in the network. The set of feasible budget vectors is denoted by $\mathcal{B}=\{\mathbf{b}\in\mathbb{R}^{m\times n}:\mathbf{b}_{si}\geqslant 0,\forall s\in\mathcal{M},\forall i\in\mathscr{N},\mathbf{1}^{\top}\mathbf{b}_{s}\leqslant C\}$ where matrix $\mathbf{b}$, represented as $(\mathbf{b}_{s},\mathbf{b}_{-s})$ denotes the joint strategy profile of all firms. Here $\mathbf{b}_{-s}$ denotes the allocation vectors of all competing firms other than $s$. Each firm is assumed to have an equal total marketing budget, $C$, which is normalized to 1. We will call this game an awareness competition game (ACG) among the firms in the rest of the paper where each player (firm) $s$ chooses their budget vector $\mathbf{b}_{s}(k)$ at every slow timescale period $k$.
Let $\mathbf{x}_{s}(t)=(x_{s1}(t),\dots,x_{sn}(t))$ denote the awareness level vector for firm $s$ at time $t$, where $x_{si}(t)\in[0,1]$ represents the awareness of node $i$ toward firm $s$. The overall awareness level of a node evolves as a linear combination of awareness gained directly through firms’ marketing efforts and indirectly through influence from neighboring nodes in the social network. To capture the temporal dynamics, awareness is modeled as a running average of updates across all preceding periods, ensuring both recent and past interactions contribute to the current state.
The basic update rule follows the framework of Bimpikis et al. (2016), which we extend to $m$ players (firms) in a two-timescale setting to separate the faster awareness dynamics from the slower strategic budget updates. Formally, the awareness of node $i$ toward firm $s$ at time $t$ of the faster timescale during the period $k$ to $k+1$ of the slower timescale is given by:
$x_{si}(t)$ $=\frac{1}{t}\left[\alpha\sum_{j\in\mathscr{N}_{i}}e_{ji}x_{sj}(t-1)+(1-\alpha)h(b_{si}(k),b_{-si}(k))\right]$
$\quad+\left(1-\frac{1}{t}\right)x_{si}(t-1)$ (1)
Here, $e_{ji}\in[0,1]$ is the $(j,i)$-th entry of the the adjacency matrix $\mathscr{E}$, that represents the influence of agent $j$ on $i$. The adjacency matrix $\mathscr{E}$ is assumed to be sub-stochastic, i.e., $\sum_{j\in\mathscr{N}_{i}}e_{ji}\leqslant 1,\;\forall i\in\mathscr{N}$, where $\mathscr{N}_{i}$ is the neighbors of $i$. The function $h(b_{si}(k),\mathbf{b}_{-si}(k))$, called the Contest Success Function (CSF), defines the probability that firm $s$ successfully influences node $i$ given the competing budget allocations. Note that the function is identical for all (firm, agent) pairs and depends only on the budget invested on node $i$ by all the firms. We will use the shorthand $h_{si}(\mathbf{b})$ to refer to $h(b_{si}(k),\mathbf{b}_{-si}(k))$ when the dependency is clear.
@@ 2.1.2 Limiting Value of Awareness Vector
The awareness for each node updates according to Equation˜1. The steady-state awareness vector for a given firm across the network is given by the following expression that follows an analysis similar to Bimpikis et al. (2016). The convergence happens between two consecutive time periods of the slow timescale $k$ as we discussed earlier. The steady-state awareness vector of firm $s$ at the end of each period $k$ is given by
@@
where $\mathbf{h}_{s}\left(\mathbf{b}_{s}(k),\mathbf{b}_{-s}(k)\right)$ is the vector $(h(b_{si}(k),\mathbf{b}_{-si}(k)),i\in\mathscr{N})$. The approximation arises due to potential convergence issues over a finite time horizon. However, for analytical tractability, we treat this relationship as an equality in the remainder of the paper. Refer to the supplementary material (Section˜A.1) for the detailed proof of this derivation, where we extend the two-firm case presented by Bimpikis et al. (2016) to the general $m$-firm setting.
@@ 2.1.3 Payoff of a firm
$u_{s}(\mathbf{b}_{s}(k),\mathbf{b}_{-s}(k))$ $=\sum_{i\in\mathscr{N}}x_{si}(k)-\sum_{i\in\mathscr{N}}b_{si}(k)$
$=\sum_{i\in\mathscr{N}}(1-\alpha)[(I-\alpha\mathscr{E})^{-1}]_{i}h_{si}(\mathbf{b}(k))-\sum_{i\in\mathscr{N}}b_{si}(k).$ (3)
@@ 2.2 Design Desiderata
A budget profile $\mathbf{b}^{*}=(\mathbf{b}_{s}^{*},\mathbf{b}_{-s}^{*})\in\mathcal{B}$ is a pure strategy Nash equilibrium (PSNE) if
$u_{s}(\mathbf{b}_{s}^{*},\mathbf{b}_{-s}^{*})\geqslant u_{s}(\mathbf{b}_{s},\mathbf{b}_{-s}^{*}),\forall\mathbf{b}_{s},\ \forall s\in\mathscr{N}.$
@@
While the competing agents reach an NE, the goal of a planner (network policymaker or regulator) is to ensure some social welfare. We consider the generalized $p$-mean welfare function, which provides a unifying framework for different welfare criteria (Barman et al., 2020).
@@
Given the $p$-mean social welfare measure $W_{p}(\mathbf{b})$ defined over all feasible budget allocation profiles $\mathbf{b}\in\mathcal{B}$, let $\mathcal{B}^{*}$ denote the set of Nash equilibria of the ACG. The Price of Anarchy (PoA) is defined as
$\text{PoA}=\frac{\textstyle\max_{\mathbf{b}\in\mathcal{B}}W_{p}(\mathbf{b})}{\textstyle\min_{\mathbf{b}\in\mathcal{B}^{*}}W_{p}(\mathbf{b})},$ (6)
@@
2.3 Existence of Nash Equilibrium
Existence of a PSNE in this context is guaranteed due to Nash’s theorem (Nash Jr, 1950). To see this, given the current game $\Gamma$, construct another game $\Gamma^{\prime}$ where the pure strategy of every firm $s\in\mathcal{M}$ is to pick one user $i\in\mathscr{N}$ on which the firm invests the entire budget of $1$. Every mixed strategy of $\Gamma^{\prime}$ is a pure strategy of $\Gamma$. Since by Nash’s theorem, a mixed strategy Nash equilibrium exists for any finite game, PSNE exists in our game. We will refer to the term PSNE as Nash equilibrium (NE) in the rest of the paper.
@@ 3 Best Response Dynamics
Even though the existence of an NE is guaranteed in a game, it is not always easy for the players to discover it and play at the equilibrium. However, there are types of games where the best response dynamics (BRD) of the players lead to an NE. BRD is an iterative method where each player sequentially updates their strategy to maximize their utility, given the current strategies of the other players. We show in this section that under certain assumptions, our awareness competition game (ACG) converges to an NE when players play BRD.
At the beginning of each slow-timescale period $k$, each firm $s\in M$ choose their budget $\mathbf{b}_{s}(k)$ while awareness levels $\mathbf{x}_{s}(t)$ evolve until time $t=T$ on that strategy profile, that is assumed to converge to a steady-state value before the next slow-timescale period $k+1$. At $k+1$, each firm updates its budget as a best response (to maximize the utility function in Equation˜4) to competitors’ previous budgets $\mathbf{b}_{-s}(k)$. Algorithm 1 illustrates the implementation of BRD for the ACG where every agent updates their budget in the direction of the gradient of their utility as given in line 4. The gradient of the utility function of firm $s$ with respect to its own budget vector $\mathbf{b}_{s}$ is defined as
$\mathbf{g}_{s}(\mathbf{b}(k),\boldsymbol{\mu}(k))$ $=\nabla_{\mathbf{b}_{s}}u_{s}(\mathbf{b}(k))+\mu_{s}(k)\in\mathbb{R}^{n\times 1}$ (7)
$\text{ where }\mu_{s}(k)$ $=\begin{cases}-\nabla_{\mathbf{b}_{s}}u_{s}(\mathbf{b}(k)),&\text{if }\mathbf{b}(k)+\gamma\mathbf{g}_{s}(\mathbf{b}(k),\boldsymbol{\mu}(k))\notin\mathcal{B}\\ 0.&\text{otherwise}\end{cases}$
@@
7:Set $\mathbf{b}^{*}\leftarrow\mathbf{b}(K)$ $\triangleright$ Nash equilibrium
8:return $\mathbf{b}^{*}$

### v2

@@ Abstract
Firms (businesses, service providers, entertainment organizations, political parties, etc.) advertise on social networks to draw people’s attention and improve their awareness of the brands of the firms. In all such cases, the competitive nature of their engagements gives rise to a game where the firms need to decide how to distribute their budget over the consumers on a network to maximize their brand’s awareness. The firms (players) therefore need to optimize how much budget they should put on the vertices (consumers) of the network so that the spread improves via direct (e.g., advertisements or free promotional offers) and indirect marketing (e.g., words-of-mouth). We propose a two-timescale model of decisions where the communication between the vertices happens in a faster timescale and the strategy update of the firms happens in a slower timescale. We show that under fairly standard conditions, the best response dynamics of the firms converge to a pure strategy Nash equilibrium. However, such equilibria can be away from a socially optimal one. We provide a characterization of the contest success functions and provide examples for the designers of such contests (e.g., regulators, social network providers, etc.) such that the Nash equilibrium becomes unique and social welfare maximizing. Our experiments show that for realistic scenarios, such contest success functions perform fairly well.
@@ 1 Introduction
In a competitive market setting, potential customers (vertex in the social graph) have varying level of information available on different firms. In this paper, we measure this information disparity a vertex has about firms through its awareness level. It captures how prominently a firm occupies an agent’s attention or belief relative to its competitors. The awareness of the customers therefore consists of two components: (a) a direct marketing impact coming from the firms through advertisements, promotional offers, etc., and (b) an indirect impact coming from recommendations from their neighbors in the social graph. The consolidated impact on the awareness therefore depends on the network centralities of the vertex w.r.t. its neighbors and from the competing budget allocated to every node by all the firms. The competing budget component is modeled through the widely used contest success function (CSF) (Skaperdas, 1996). Our analysis considers that the firms play best response to each other’s committed budgets to reach equilibrium, and analyzes the welfare at such equilibria. The CSF is viewed as a design choice made by a planner who sets the rules governing how firms’ marketing efforts translate into customer awareness, with the goal of shaping firms’ incentives so that their strategic behavior leads to socially desirable outcomes. This perspective is consistent with real-world advertising markets, where platform owners such as Google or Meta determine how impressions are allocated among competing advertisers through platform-specific rules, based on bids or budgets.
@@ 1.1 Our Contributions
The contributions of this paper are summarized as follows.
  • We introduce a fast and slow timescale model of awareness spread in a social network where the firms update their budget strategies on the slow timescale and the individuals update their awareness on the faster timescale (Section˜2.1).
@@
  • On the slower timescale, we show that the budget strategy of the firms converges to a Nash equilibrium (NE) (Theorem˜1). However, the social welfare at the Nash equilibrium may not be optimal (˜1).
@@
For better readability, some details of the proofs and experiments have been deferred to the supplementary material.
@@ 1.2 Related Work
Modeling and structural guarantees of competitive influence.
A significant line of research has focused on formalizing competitive diffusion processes and analyzing the properties of equilibria. Foundational work by Bharathi et al. (2007) study the diffusion of competing innovations in social networks, proposing a tractable game-theoretic model, deriving a (11/e)(1-1/e)-approximation algorithm for best responses, and bounding the price of competition. Early works (Alon et al., 2010; Goyal and Kearns, 2012) formalize the game theoretic competitive influence maximizing model and emphasize the role of topology in the existence of Nash equilibrium and influence amplification. Zuo et al. (2022) introduce the Online Competitive Influence Maximization (OCIM) problem, and design bandit-based algorithms with sublinear regret. Tzoumas et al. (2012) models the diffusion process of CIM in a simultaneous two-player non-cooperative game as a linear threshold model. Fazeli and Jadbabaie (2012a) and Fazeli and Jadbabaie (2012b); Fazeli et al. (2016) develop a series of game-theoretic models for competitive product diffusion in social networks, where firms simultaneously allocate fixed budgets between seeding strategies and product quality, aiming to maximize the lower bound of their product adoptions.
@@ 2 Preliminaries
We consider a market with $m$ competing firms denoted by $\mathcal{M}=\{1,2,\ldots,m\}$, each having an equal and fixed marketing budget that they strategically allocate across users in a social network. Firms compete to attract the attention of these users by distributing their budgets to maximize brand visibility and influence. We consider awareness as a cardinal metric of a user’s attention towards a particular firm, which lies within $[0,1]$. This marketing competition is modeled as a finite non-cooperative game on a social network $\mathscr{G}=(\mathscr{N},\mathscr{E})$, where $\mathscr{N}=\{1,2,\ldots,n\}$ denotes the set of nodes/users in the network and $\mathscr{E}=(e_{ij})_{i,j\in\mathscr{N}}$ represents their influence structure through an adjacency matrix, with $e_{ii}=0$ for all $i\in\mathscr{N}$. Each firm aims to maximize the overall awareness of its brand across the network, subject to its budget constraints.
@@ 2.1 A Two-Timescale Communication Model
We model the interaction between firms’ strategic decisions and the evolution of customer awareness using a two-timescale framework.
Slow timescale (strategy updates): The slower timescale, indexed by $k=\{0,1,\ldots,\infty\}$, governs firms’ strategic decisions in the form of budget allocations. At the beginning of each period $k$, firms allocate their budgets across users through direct communication with a probability $(1-\alpha)$ where $\alpha\in(0,1)$. These allocations determine the contest outcome via the CSF and remain fixed throughout the entire period.
Fast timescale (awareness dynamics): Within each period $k$, the faster timescale variable $t=\{0,1,\ldots,T\}$ captures how awareness propagates and stabilizes among users, where at each timestep $t$, peer-to-peer communication among nodes occurs with probability $\alpha$. This process occurs at a much finer temporal resolution than firms’ strategy updates, reflecting the fact that user interactions, information sharing, and content consumption on social platforms take place frequently and continuously. We assume that $T$ is sufficiently large so that the fast-timescale awareness dynamics converge to a steady state before the next strategy update at period $k+1$. This separation of timescales is consistent with real-world advertising settings, where communication and information diffusion among users typically occur far more frequently than firms revise their advertising budgets and strategies. From a technical perspective, this assumption also enables a tractable analysis by allowing the convergence of awareness on the fast timescale and the budget updates on the slow timescale to be studied independently. While scenarios in which the fast-timescale dynamics do not fully converge within a single period can be explored through simulations, their analytical treatment would be considerably more involved and is beyond the scope of this work. Accordingly, throughout the paper we overload $t$ to denote the fast-timescale discrete variable running from 0 to $T$ within each period $k$.
@@ 2.1.1 Awareness levels of each firm
Each firm $s\in M$ chooses a strategy vector $\mathbf{b}_{s}=(b_{s,1},\dots,b_{s,n})$ representing the allocation of its marketing budget across the $n$ nodes in the network. The set of feasible budget vectors is denoted by $\mathcal{B}=\{\mathbf{b}\in\mathbb{R}^{m\times n}:b_{s,i}\geqslant 0,\forall s\in\mathcal{M},\forall i\in\mathscr{N},\mathbf{1}^{\top}\mathbf{b}_{s}\leqslant C\}$ where matrix $\mathbf{b}$, represented as $(\mathbf{b}_{s},\mathbf{b}_{-s})$ denotes the joint strategy profile of all firms. Here $\mathbf{b}_{-s}$ denotes the allocation vectors of all competing firms other than $s$. Each firm is assumed to have an equal total marketing budget, $C$, which is normalized to 1. We will call this game an awareness competition game (ACG) among the firms in the rest of the paper where each player (firm) $s$ chooses their budget vector $\mathbf{b}_{s}(k)$ at every slow timescale period $k$.
Let $\mathbf{x}_{s}(t)=(x_{s,1}(t),\dots,x_{s,n}(t))$ denote the awareness level vector for firm $s$ at time $t$, where $x_{s,i}(t)\in[0,1]$ represents the awareness of node $i$ toward firm $s$. The overall awareness level of a node evolves as a linear combination of awareness gained directly through firms’ marketing efforts and indirectly through influence from neighboring nodes in the social network. To capture the temporal dynamics, awareness is modeled as a running average of updates across all preceding periods, ensuring both recent and past interactions contribute to the current state.
The basic update rule follows the framework of Bimpikis et al. (2016), which we extend to $m$ firms in a two-timescale setting to separate the faster awareness dynamics from the slower strategic budget updates. Formally, the awareness of node $i$ toward firm $s$ at time $t$ of the faster timescale during the period $k$ to $k+1$ of the slower timescale is given by:
$x_{s,i}(t)$ $=\frac{1}{t}\left[\alpha\sum_{j\in\mathscr{N}_{i}}e_{ji}x_{s,j}(t-1)+(1-\alpha)h(b_{s,i}(k),\mathbf{b}_{-s,i}(k))\right]$
$\quad+\left(1-\frac{1}{t}\right)x_{s,i}(t-1)$ (1)
Here, $e_{ji}\in[0,1]$ is the $(j,i)$-th entry of the the adjacency matrix $\mathscr{E}$, that represents the influence of node $j$ on $i$. The adjacency matrix $\mathscr{E}$ is assumed to be sub-stochastic, i.e., $\sum_{j\in\mathscr{N}_{i}}e_{ji}\leqslant 1,\;\forall i\in\mathscr{N}$, where $\mathscr{N}_{i}$ is the neighbors of $i$. The function $h(b_{s,i}(k),\mathbf{b}_{-s,i}(k))$, called the Contest Success Function (CSF), defines the probability that firm $s$ successfully influences node $i$ given the competing budget allocations. Here, $b_{s,i}$ denotes the scalar budget allocated by firm $s$ to node $i$, while $\mathbf{b}_{-s,i}$ denotes the vector of budgets allocated to node $i$ by all firms other than $s$. Note that the function is identical for all (firm, node) pairs and depends only on the budget invested on node $i$ by all the firms. We will use the shorthand $h_{s,i}(\mathbf{b})$ to refer to $h(b_{s,i}(k),\mathbf{b}_{-s,i}(k))$ when the dependency is clear.
@@ 2.1.2 Limiting Value of Awareness Vector
The awareness for each node updates according to Equation˜1. The steady-state awareness vector for a given firm across the network is given by the following expression that follows an analysis similar to Bimpikis et al. (2016). The convergence happens between two consecutive periods of the slow timescale $k$ as we discussed earlier. The steady-state awareness vector of firm $s$ at the end of each period $k$ is given by
@@
where $\mathbf{h}_{s}\left(\mathbf{b}_{s}(k),\mathbf{b}_{-s}(k)\right)$ is the vector $(h(b_{s,i}(k),\mathbf{b}_{-s,i}(k)),i\in\mathscr{N})$. The approximation arises due to potential convergence issues over a finite time horizon. However, for analytical tractability, we treat this relationship as an equality in the remainder of the paper. Refer to the supplementary material (Section˜A.1) for the detailed proof of this derivation, where we extend the two-firm case presented by Bimpikis et al. (2016) to the general $m$-firm setting.
@@ 2.1.3 Payoff of a firm
$u_{s}(\mathbf{b}_{s}(k),\mathbf{b}_{-s}(k))$ $=\sum_{i\in\mathscr{N}}x_{s,i}(k)-\sum_{i\in\mathscr{N}}b_{s,i}(k)$
$=\sum_{i\in\mathscr{N}}(1-\alpha)[(I-\alpha\mathscr{E})^{-1}]_{i}h_{s,i}(\mathbf{b}(k))-\sum_{i\in\mathscr{N}}b_{s,i}(k).$ (3)
@@ 2.2 Design Desiderata
A budget profile $\mathbf{b}^{\star}=(\mathbf{b}_{s}^{\star},\mathbf{b}_{-s}^{\star})\in\mathcal{B}$ is a pure strategy Nash equilibrium (PSNE) if
$u_{s}(\mathbf{b}_{s}^{\star},\mathbf{b}_{-s}^{\star})\geqslant u_{s}(\mathbf{b}_{s},\mathbf{b}_{-s}^{\star}),\forall\;\mathbf{b}_{s},\ \forall s\in\mathscr{N}.$
@@
While the competing firms reach an NE, the goal of a planner (network policymaker or regulator) is to ensure some social welfare. We consider the generalized $p$-mean welfare function, which provides a unifying framework for different welfare criteria (Barman et al., 2020).
@@
Given the $p$-mean social welfare measure $W_{p}(\mathbf{b})$ defined over all feasible budget allocation profiles $\mathbf{b}\in\mathcal{B}$, let $\mathcal{B}^{\star}$ denote the set of Nash equilibria of the ACG. The Price of Anarchy (PoA) is defined as
$\text{PoA}=\frac{\textstyle\max_{\mathbf{b}\in\mathcal{B}}W_{p}(\mathbf{b})}{\textstyle\min_{\mathbf{b}\in\mathcal{B}^{\star}}W_{p}(\mathbf{b})},$ (6)
@@ 3 Best Response Dynamics
Even though the existence of an NE is guaranteed in a game, it is not always easy for the players to discover it and play at the equilibrium. However, there are types of games where the best response dynamics (BRD) of the players lead to an NE. BRD is an iterative method where each player sequentially updates their strategy to maximize their utility, given the current strategies of the other players. We show in this section that under certain assumptions, our awareness competition game (ACG) converges to an NE when players (firms) play BRD.
At the beginning of each slow-timescale period $k$, each firm $s\in M$ choose their budget $\mathbf{b}_{s}(k)$ while awareness levels $\mathbf{x}_{s}(t)$ evolve until time $t=T$ on that strategy profile, that is assumed to converge to a steady-state value before the next slow-timescale period $k+1$. At $k+1$, each firm updates its budget as a best response (to maximize the utility function in Equation˜4) to competitors’ previous budgets $\mathbf{b}_{-s}(k)$. Algorithm 1 illustrates the implementation of BRD for the ACG where every firm updates their budget in the direction of the gradient of their utility as given in line 4. The gradient of the utility function of firm $s$ with respect to its own budget vector $\mathbf{b}_{s}$ is defined as $\mathbf{g}_{s}(\mathbf{b}(k),\boldsymbol{\mu}(k))\in\mathbb{R}^{n\times 1}$ where its $i^{th}$ component is given by
$g_{s,i}(\mathbf{b}(k),\boldsymbol{\mu}_{s}(k))$ $=\frac{\partial u_{s}(\mathbf{b}(k))}{\partial b_{s,i}}+\mu_{s,i}(k),$
$\text{ where }\mu_{s,i}(k)$ $=\begin{cases}-\dfrac{\partial u_{s}}{\partial b_{s,i}}(\mathbf{b}(k)),&\text{if }b_{s,i}(k)=0\text{ and }\dfrac{\partial u_{s}}{\partial b_{s,i}}(\mathbf{b}(k))<0,\\[10.00002pt] -\dfrac{\partial u_{s}}{\partial b_{s,i}}(\mathbf{b}(k)),&\text{if }\sum_{i}b_{s,i}(k)=C\text{ and }\dfrac{\partial u_{s}}{\partial b_{s,i}}(\mathbf{b}(k))>0,\\[10.00002pt] 0,&\text{otherwise}.\end{cases}$
Hence, the gradient vector can be expressed as:
$\mathbf{g}_{s}(\mathbf{b}(k),\boldsymbol{\mu}(k))=\nabla_{\mathbf{b}_{s}}u_{s}(\mathbf{b}(k))+\boldsymbol{\mu}_{s}(k)$ (7)
Intuition. The role of the correction term $\boldsymbol{\mu}_{s}(k)$ is to remove only those components of the gradient that would drive the update outside the feasible set. Recall that the feasible set for each firm $s$ is $\mathcal{B}_{s}=\{\mathbf{b}_{s}\in\mathbb{R}_{+}^{n}:\sum_{i}b_{s,i}\leq C\}$. A strategy $\mathbf{b}_{s}$ lies on the boundary of $\mathcal{B}_{s}$ if either $b_{s,i}=0$ for some $i$ or $\sum_{i}b_{s,i}=C$. At such boundary points, only those components of the gradient are canceled that point outside $\mathcal{B}_{s}$, while any feasible budget redistribution direction at the boundary that increases utility is preserved in the update direction $\mathbf{g}_{s}(\mathbf{b}(k),\boldsymbol{\mu}(k))$.
@@
7:Set $\mathbf{b}^{\star}\leftarrow\mathbf{b}(K)$ $\triangleright$ Nash equilibrium
8:return $\mathbf{b}^{\star}$
The following lemma characterizes the stationary points of the proposed best-response dynamics and establishes their relationship to pure-strategy Nash equilibria.
Lemma 1.
A strategy profile $\mathbf{b}^{\star}$ satisfies
$\mathbf{g}_{s}(\mathbf{b}^{\star},\boldsymbol{\mu}_{s}^{\star})=\mathbf{0}\quad\text{for all }s\in\mathcal{M}$
if and only if $\mathbf{b}^{\star}$ is a pure-strategy Nash equilibrium according to Definition˜1.
Proof.
Suppose first that
$\mathbf{g}_{s}(\mathbf{b}^{\star},\boldsymbol{\mu}_{s}^{\star})=\mathbf{0}\quad\text{for all }s\in\mathcal{M}.$
Assume, for the sake of contradiction, that $\mathbf{b}^{\star}$ is not a Nash equilibrium. Then there exists a firm $s$ and a feasible unilateral deviation $\tilde{\mathbf{b}}_{s}\in\mathcal{B}_{s}$ such that
$u_{s}(\tilde{\mathbf{b}}_{s},\mathbf{b}_{-s}^{\star})>u_{s}(\mathbf{b}_{s}^{\star},\mathbf{b}_{-s}^{\star}).$
By construction of the update rule, this implies that the gradient of $u_{s}$ at $\mathbf{b}^{\star}$ admits a feasible improving component. Consequently, the correction term $\boldsymbol{\mu}_{s}^{\star}$ does not cancel this component of the gradient, and the feasibility-corrected update direction $\mathbf{g}_{s}(\mathbf{b}^{\star},\boldsymbol{\mu}_{s}^{\star})$ must be nonzero, i.e. $\mathbf{g}_{s}(\mathbf{b}^{\star},\boldsymbol{\mu}_{s}^{\star})\neq\mathbf{0}$, which contradicts the assumption. Therefore, no firm admits a profitable feasible deviation at $\mathbf{b}^{\star}$, and $\mathbf{b}^{\star}$ is a pure-strategy Nash equilibrium according to Definition˜1. ∎
The above lemma clarifies that the proposed best-response dynamics cannot become stationary at a non-equilibrium profile as the dynamics do not converge to a trivial zero-gradient state.